OpenStudy (anonymous):

Help? Find all solutions in the interval [0, 2pi). sec^2 x – 2 = tan^2 x

OpenStudy (tkhunny):

This is not a calculation problem. This is a paying attention problem. Really Big Hint: \(\sec^{2}(x) - \tan^{2}(x) = 1\) You just have to see those easy ones and reach out and grab them!

OpenStudy (anonymous):

are you sure this is the question?

OpenStudy (dumbcow):

rearranging gives \[\sec^{2} x - \tan^{2} x = 2\] which is false .... thus No solution

OpenStudy (anonymous):

if so, there is no solution subtract \(\tan^2(x)\) from both sides, get \[\sec^2(x)-\tan^2(x)-2=0\]\[1-2=0\] \[-1=0\] no solution

OpenStudy (anonymous):

But wouldnt sec^2 x - 2 tan^2 x = 2?

OpenStudy (anonymous):

Oooh okay, nevermind. I understand. Thank you :o)

OpenStudy (anonymous):

you only have \(\tan^2(x)\) on the right, yes?

OpenStudy (anonymous):

Yeah, I got confused with the -2 on the left.